OpenStudy (kittiwitti1):

At what Celsius temperature will the numerical reading on the Fahrenheit thermometer be 51 ∘ less than that on the Celsius thermometer? Express your answer using two significant figures. I keep getting the wrong answer...

OpenStudy (kittiwitti1):

F to C: (C-32)/1.8 Problem interpretation: F = C-51

OpenStudy (kittiwitti1):

Therefore: C-32/1.8 = C-51?

OpenStudy (mathstudent55):

C-32/1.8 is not degrees F.

OpenStudy (kittiwitti1):

Oh... :(

OpenStudy (mathstudent55):

F = 1.8C + 32

OpenStudy (kittiwitti1):

I meant "F-32/1.8" my bad

OpenStudy (mathstudent55):

C = (F - 32)/1.8

OpenStudy (kittiwitti1):

Yeah. That... ugh

OpenStudy (mathstudent55):

You must put parentheses around F - 32.

OpenStudy (kittiwitti1):

Alright. How would I get only C in the equation, though?

OpenStudy (mathstudent55):

You do a substitution. Start with: F = C - 51

OpenStudy (mathstudent55):

F = 1.8C + 32, so where you see F, replace F with what F is equal to, which is 1.8C + 32. 1.8C + 32 = C - 51

OpenStudy (kittiwitti1):

Oh... is this going to turn into an equivalence problem?

OpenStudy (kittiwitti1):

... yup

OpenStudy (kittiwitti1):

Okay, so then I would end up with 0.8C = (-51-32)?

OpenStudy (mathstudent55):

correct

OpenStudy (kittiwitti1):

Okay... I got this: 0.8C= -83 C = -103.75 degrees.

OpenStudy (mathstudent55):

Correct.

OpenStudy (kittiwitti1):

And to cut down to two SF would make it ... I'm not sure how to do that actually

OpenStudy (kittiwitti1):

-103 rounds to -100, but that's three SF

OpenStudy (mathstudent55):

Now let's check. Let's use our formula above to change -103.75 deg C to F F = 1.8C + 32 F = 1.8(-103.75) + 32 F = -186.75 + 32 F = -154.75 Sure enough, the number -154.75 51 less than -103.75

OpenStudy (kittiwitti1):

Yeah, but I don't know how to round sig figs... it wants two, but the only rounding I get is three sig figs. e-e

OpenStudy (mathstudent55):

I see. You need 2 significant figures.

OpenStudy (kittiwitti1):

Although I'm really glad you helped me through the other parts. Thanks for that

OpenStudy (kittiwitti1):

I used a calculator and it gave me:\[1.0\times10^{2}\]

OpenStudy (mathstudent55):

I think you can use scientific notation. -103.75 in scientific notation is -1.0375 * 10^2 Now write -1.0375 with 2 significant figures.

OpenStudy (kittiwitti1):

I thought 1.00 was the proper scientific notation?

OpenStudy (mathstudent55):

Correct for the significant figure part, but remember it is negative.

OpenStudy (mathstudent55):

1.00 has 3 sig figures. You must have 2, not 3, so use 1.0

OpenStudy (kittiwitti1):

Ah okay

OpenStudy (kittiwitti1):

... it said I got it wrong...

OpenStudy (kittiwitti1):

The last two were my tries at the answer we got. http://prntscr.com/cfh3sx

OpenStudy (kittiwitti1):

OH

OpenStudy (kittiwitti1):

ARGH

OpenStudy (mathstudent55):

It says "check your signs." I mentioned above the answer is negative, so you need a negative sign before the 1.0

OpenStudy (mathstudent55):

-1.0 * 10^2

OpenStudy (mathstudent55):

Not so loud...

OpenStudy (kittiwitti1):

Ahem. * (*silent screaming*)

OpenStudy (kittiwitti1):

I'm sorry, I'm an idiot

OpenStudy (mathstudent55):

If you were an idiot, you'd have given up. The fact that you're still at it until you understand it and get it right proves you are not.

OpenStudy (kittiwitti1):

At what Celsius temperature will the numerical reading on the Fahrenheit thermometer be 3-times more that on the Celsius thermometer?

OpenStudy (kittiwitti1):

Aw, thank you.

OpenStudy (mathstudent55):

Set it up just like you did the first problem. First, write an equation with F and C that is just a translation of the sentence.

OpenStudy (mathstudent55):

the numerical reading on the Fahrenheit thermometer be 3-times more that on the Celsius thermometer

OpenStudy (kittiwitti1):

So another substitution?

OpenStudy (mathstudent55):

Yes, but we are still working on the step before the substitution.

OpenStudy (mathstudent55):

What is "3-times more that on the Celsius thermometer", using C for the Celsius reading?

OpenStudy (kittiwitti1):

3x more is just times 3 right?

OpenStudy (mathstudent55):

3 times the Celsius reading 3 times C

OpenStudy (kittiwitti1):

F = 3C and then combine that with the standard F conversion formula?

OpenStudy (mathstudent55):

Ok. Now we use the conversion formula and we substitute.

OpenStudy (kittiwitti1):

So... 1.8C+32 = 3C

OpenStudy (mathstudent55):

I just want to explain something before we substitute.

OpenStudy (kittiwitti1):

Oh. Yes?

OpenStudy (mathstudent55):

There are two conversion formulas for F to C and for C to F. The are: C = (F - 32)/1.8 to convert from F to C F = 1.8C + 32 to convert from C to F How do we know which one to choose for the substitution step?

OpenStudy (kittiwitti1):

The F, because F=3C

OpenStudy (kittiwitti1):

F to C, I mean

OpenStudy (kittiwitti1):

GAH.

OpenStudy (kittiwitti1):

Oh, and I got this one right. Off by a couple digits but still correct

OpenStudy (mathstudent55):

We go back to the question. The question is asking for the Celsius temperature. We need to solve for C, There is a solution for F also, but we were not asked to find it. Since we are solving only for C, we need to get rid of F and end up with only C in the equation. That is why we look for a conversion that has F = some expression in C. In this case, it is F = 1.8C + 32 After we substitute the conversion, we end up with an equation with just the variable C, so we can solve for C.

OpenStudy (kittiwitti1):

I got the answer right :)

OpenStudy (kittiwitti1):

At what Celsius temperature will the numerical reading on the Fahrenheit thermometer be 1/8 that on the Celsius thermometer?

OpenStudy (kittiwitti1):

Uh.. is this F = 1/8 C or F = 8 C?

OpenStudy (mathstudent55):

One again, we now have F = 3C To get rid of F, we use F = 1.8C + 32 1.8C + 32 = 3C

OpenStudy (mathstudent55):

F = (1/8)C

OpenStudy (mathstudent55):

and now substitute the same way as before.

OpenStudy (kittiwitti1):

Alright, thank you :)

OpenStudy (kittiwitti1):

Mom's insulting me for being on here to get homework help...

OpenStudy (mathstudent55):

Listen to your mom. Bye.

OpenStudy (kittiwitti1):

... I'm trying to work on my homework, though.